# Continuous time asymptotic representations for adaptive experiments

### v1

@@ 2.3.2. Allocation processes vs policy rules
However, $\pi(\cdot)\equiv\pi_{1}(\cdot)$ cannot, in general, be interpreted as a valid policy rule, since it need not be measurable as a random process indexed by $t$. Indeed, in many continuous-time optimal control problems, the optimal policy does not belong to a separable space, and is therefore not measurable. [^4] The measurability of sample paths—often referred to as strong or Bochner measurability—differs from the weaker notion of coordinate-wise measurability. The Pettis measurability theorem states that a stochastic process is Bochner measurable if and only if its sample paths lie in a separable subspace with probability one. See adusumilli2023optimal for further discussion. For this reason, it is also generally impossible to define weak convergence for a sequence of policies $\pi_{n,\lfloor nt\rfloor}$, as such sequences are typically not asymptotically equicontinuous.
@@ 3. Asymptotic equivalence of experiments
We now establish the asymptotic equivalence between the original sequence of adaptive experiments and the limit experiment. This equivalence follows from two key results. The first, previously demonstrated in adusumilli2025optimal, states that the likelihood ratio processes in the original experiment converge uniformly to their counterparts in the limit experiment. We restate this result here for completeness. The second result, which is novel to this article, asserts that any sequence of score and allocation processes, $\{x_{n,a}(\cdot),q_{n,a}(\cdot)\}_{a}$, admits a corresponding representation in the limit experiment.
@@ 3.1. Convergence of likelihood ratio processes
In adusumilli2025optimal, this author showed that under Assumption 1,
@@
In other words, the likelihood ratio processes under the original experiment converge uniformly to those under the limit experiment for each $\bm{h}$. Following blackwell1953equivalent and le1986asymptotic, this implies that the two experiments are equivalent, in that the posteriors converge uniformly over all possible values of $\gamma_{1},\gamma_{0}$.
By itself, (3.3) does not make any reference to an allocation process. To show that a risk function in the original experiment admits a corresponding representation or lower bound in the limit experiment, we need to go further and match the allocation processes as well. In particular, we need to show that for any set of (asymptotically) sufficient statistics, $\{x_{n,a}(\cdot),q_{n,a}(\cdot)\}_{a}$, there exists a corresponding $\{x_{a}(\cdot),q_{a}(\cdot)\}_{a}$ in the limit experiment such that, when $\bm{h}=(0,0)$, the distribution of the former converges to that of the latter as $n\to\infty$. This is the key result that we prove next.
@@ 3.2. The main result
This ensures that $\{z_{n,a}(\cdot),q_{n,a}(\cdot)\}_{a}$ converges to some weak limit $\{z_{a}(\cdot),q_{a}(\cdot)\}_{a}$. Moreover, the measurability of the events $\{q_{n,1}(t)\leq\gamma_{1},q_{n,0}(t)\leq\gamma_{0}\}$ with respect to $\mathcal{G}_{n,\gamma_{0},\gamma_{1}}:=\mathcal{F}_{n,\gamma_{1}}^{(1)}\lor\mathcal{F}_{n,\gamma_{0}}^{(0)}\lor\sigma(U)$ suggests that $q_{a}(\cdot)$ can be constructed to inherit the appropriate adaptedness properties required by Definition 1. The construction requires some care and utilizes some results from the theory of stable convergence (hausler2015stable); it is perhaps the most intricate part of the proof. Setting $x_{n,a}(\cdot):=z_{n,a}(q_{n,a}(\cdot))$ and $x_{a}(\cdot):=z_{a}(q_{a}(\cdot))$ then gives the desired result.
@@ 4.1. Bayes risk
Let $m_{\bm{\theta}}(\cdot)$ denote a given prior over $\bm{\theta}$. This induces a local prior, $m(\cdot)$, over $\bm{h}$ through the transformation $\bm{h}=\sqrt{n}(\bm{\theta}-\bm{\theta}_{0})$. We consider an asymptotic regime wherein $m(\cdot)$ is assumed to be independent of $n$. The influence of the prior thus remains asymptotically non-negligible and the Bernstein-von Mises theorem does not apply. As discussed in adusumilli2025optimal, local priors offer a better approximation to finite-sample behavior because their influence does not diminish with sample size.
@@ 5. Application 2: Equivalence of in-sample regret
Let $\mu_{n,a}(\bm{h}):=\mathbb{E}_{n,\bm{h}}[Y_{i}^{(a)}]$ denote the average reward corresponding to treatment $a$ when the local parameter is $\bm{h}$. Following hirano2023asymptotic and adusumilli2025optimal, the reference parameter is chosen such that $\mu_{n,a}(\bm{0})=0$. The frequentist regret of any $\pi_{n}\equiv\{\pi_{n,j}\}_{j}$ is given by
@@
Previous work by this author (adusumilli2025optimal) characterized the optimal Bayes and minimax risks in this setting, and showed that they can be attained by a sequence of policy rules that depend on just $\left\{x_{a}(t),q_{a}(t)\right\}_{a}$, i.e., the past values of these variables are not relevant. Theorem 3 is more general, in that it applies to arbitrary sequences of policy rules, but it makes use of a larger information set that includes the entire sample paths of $\{x_{a}(s);s\leq t\}_{a}$ until time $t$.
@@ 6.1. E-processes in multi-treatment adaptive experiments
Definition 2 generalizes the the usual notion of an e-process to a multi-indexed super-martingale, where the indices are the treatment allocation proportions. In this framework, the value of $\varepsilon_{n}(\cdot)$ depends on the trajectory of the empirical allocation process, and the super-martingale property holds across all possible empirical allocation processes, i.e., all possible adaptive experiments. A central feature of the e-process is that it is algorithm-free: it remains a valid supermartingale at any time point of any adaptive experiment.
@@ 6.2. GRO, mGRO and REGROW
Following ramdas2023game and grunwald2024safe, a common approach to evaluating e-processes in stopping-time experiments is through the Growth Rate Optimality (GRO) criterion. GRO assesses the quality of an e-process $\varepsilon_{n}(\cdot)$ based on its expected log-growth under an alternative hypothesis $\bm{h}\in\mathcal{H}_{1}$. We extend this criterion to the multi-treatment adaptive setting by defining the GRO score
@@
An alternative criterion, following grunwald2024safe, is the REGROW (RElative GRowth Optimality in Worst case) score:
@@ 6.3. Local asymptotics and e-processes in the limit experiment
We then define an equivalent notion of an e-process in the limit experiment. Recall that the limit experiment is characterized by Gaussian process signals $z_{a}(q):=I_{a}^{1/2}h^{(a)}q+W_{a}(q)$. The e-process in the limit experiment is a $\mathcal{G}_{q_{1},q_{0}}$-adapted stochastic process designed for testing $H_{0}:\bm{h}\in\mathcal{H}_{0}$ versus $H_{1}:\bm{h}\in\mathcal{H}_{1}$. The formal definition is given below. First, a bit of terminology: let $\bm{\mathcal{Q}}$ denote the collection of allocation processes in the limit experiment that are weak limits of some sequence of empirical allocation processes $\{q_{n,a}(\cdot)\}_{a}$.
@@
Relative to Definition 2, the above definition restricts the set of allocation processes to $\bm{\mathcal{Q}}$. The rationale behind this is mainly technical: it allows us to avoid dealing with $\{q_{a}(\cdot)\}_{a}$ that are not weak limit points of sequences of empirical allocation processes in the actual experiment. Whether $\bm{\mathcal{Q}}$, in fact, includes all possible allocation processes is currently unknown (to this author).
@@ 6.5. Applying the representation theorems
In what follows, we simplify matters by assuming that the null hypothesis $\mathcal{H}_{0}$ is a singleton, consisting solely of the reference parameter $\bm{\theta}_{0}$. [^7] When there is only adaptive stopping, compound nulls are typically addressed in anytime-valid inference using the method of reverse information projection (see ramdas2023game, for a survey). Extending this approach to the adaptive sampling setting is more involved and left for future work. See, however, Appendix E for the simpler case of testing parameters corresponding to a single treatment arm. Theorem 4 and Corollaries 3, 4 establish asymptotic upper bounds on the GRO, mGRO and REGROW criteria. These bounds are obtained by optimizing the respective criteria within the limit experiment.
@@ 6.5.2. REGROW optimality
grunwald2024safe show that the REGROW-optimal e-value, $\bar{\varepsilon}_{q_{1}(t),q_{0}(t)}^{*}$, coincides with the mGRO-optimal e-value for a specific, least-favorable weighting function $w_{q_{1}(t),q_{0}(t)}^{*}(\cdot)$. The authors also show that this weighting function is obtained as the solution to
@@
grunwald2024safe further demonstrate that the optimized value of the objective in (6.5) provides an upper bound on the REGROW criterion $\mathcal{R}\left(\varepsilon;\{q_{a}(t)\}_{a}\right)$:
@@
Crucially, the least-favorable distribution $w_{(\cdot)}^{*}$ depends on both the chosen allocation process and the structure of the alternative hypothesis set $\mathcal{H}_{1}$. It is essential that $\mathcal{H}_{1}$ be compact—without it, the channel capacity (and hence the REGROW value) becomes infinite. Because of the dependence on the allocation process, a globally optimal REGROW e-process does not exist. In practice, one must fix a specific allocation process $\{q_{a}(t)\}_{a}$ and an alternative region $\mathcal{H}_{1}$, from which an optimal weighting function $w_{(\cdot)}^{*}$ can be derived, e.g., by employing the Blahut-Arimoto algorithm (see, arimoto1972algorithm; blahut1972computation). The resulting e-process is thus locally REGROW optimal relative to the chosen design and alternative set (it would also be globally mGRO optimal relative to the least favorable distribution $w^{*}(\cdot)$). Appendix E provides an example of such a locally REGROW optimal e-process, constructed to be optimal against fixed values of $\{q_{a}\}_{a}$.
@@ 7. Conclusion
This article introduces a continuous-time formalism for analyzing fully adaptive experiments. The formalism is based on the notion of allocation processes, also introduced in this article. We show that any empirical allocation process, as induced by some policy rule, converges weakly to a corresponding allocation process in a limit experiment governed by Gaussian diffusions. The various applications, ranging from point-estimation to anytime-valid inference illustrate the utility and generality of this framework.
Beyond these applications, the continuous-time formulation offers a powerful tool for addressing design problems in adaptive experimentation. Though not based on the results reported in this article, prior work by this author has applied the continuous-time framework to derive optimal algorithms for bandit experiments and costly sampling problems (adusumilli2025optimal; adusumilli2025sample). The current results provide an easy-to-use and generalizable template for transferring optimal designs from the limit experiment back to the finite-sample setting. Looking ahead, we expect this approach to be broadly useful in addressing a range of open questions in adaptive experimentation—e.g., in deriving optimal strategies for best-arm identification with multiple treatments, or in designing adaptive experiments in strategic environments involving interactions between an experimenter and a regulator.
@@ A.1. Proof of Lemma 1
Fix any $t_{1},t_{2}\in[0,1]$ such that $t_{2}\geq t_{1}$, and define $\tau_{1}:=q_{1}(t_{1}),\tau_{2}:=q_{1}(t_{2})$. Also, for each $\gamma_{1}\geq 0$, define $\mathcal{H}_{\gamma_{1}}:=\mathcal{G}_{\gamma_{1},1}$. By the definition of the allocation process, the event $\{q_{1}(t)\leq\gamma_{1}\}$ is $\mathcal{H}_{\gamma_{1}}$ measurable, under any given $t$. Hence, $\tau_{1},\tau_{2}$ are both $\{\mathcal{H}_{\gamma_{1}}\}_{\gamma_{1}\geq 0}$-adapted stopping times. It is easily verified from the definition of $\mathcal{G}_{\gamma_{1},\gamma_{0}}$ that $z_{1}(\gamma_{1})$ is a Wiener process with respect to $\mathcal{H}_{\gamma_{1}}$. Since $\tau_{2}\geq\tau_{1}$ almost surely (a.s.,) due to the almost sure monotonicity of $q(\cdot)$, it follows by the optional sampling theorem that $\mathbb{E}\left[z(\tau_{2})|\mathcal{H}_{\tau_{1}}\right]=z(\tau_{1})$ a.s. In other words, $\mathbb{E}\left[z_{1}(q_{1}(t_{2}))|\mathcal{H}_{q_{1}(t_{1})}\right]=z_{1}(q_{1}(t_{1}))$ a.s., i.e., $\mathbb{E}\left[x_{1}(t_{2})|\mathcal{H}_{q_{1}(t_{1})}\right]=x_{1}(t_{1})$. Since $\mathcal{I}_{t}\equiv\mathcal{G}_{q_{1}(t),q_{0}(t)}\subseteq\mathcal{G}_{q_{1}(t),1}\equiv\mathcal{H}_{q_{1}(t)}$ for any $t$, the tower property of conditional expectations then implies $\mathbb{E}\left[x_{1}(t_{2})|\mathcal{I}_{t_{1}}\right]=x_{1}(t_{1})$ a.s. But $t_{1},t_{2}$ were arbitrary, so the claim follows.

### v2

@@ 2.3.2. Allocation processes vs policy rules
However, $\pi(\cdot)\equiv\pi_{1}(\cdot)$ cannot, in general, be interpreted as a valid policy rule, since it need not be measurable as a random process indexed by $t$. Indeed, in many continuous-time optimal control problems, the optimal policy does not belong to a separable space, and is therefore not measurable. [^5] The measurability of sample paths—often referred to as strong or Bochner measurability—differs from the weaker notion of coordinate-wise measurability. The Pettis measurability theorem states that a stochastic process is Bochner measurable if and only if its sample paths lie in a separable subspace with probability one. See Adusumilli (2023, Section 5.1.1) for further discussion. For this reason, it is also generally impossible to define weak convergence for a sequence of policies $\pi_{n,\lfloor nt\rfloor}$, as such sequences are typically not asymptotically equicontinuous.
@@ 3. Asymptotic equivalence of experiments
We now establish the asymptotic equivalence between the original sequence of adaptive experiments and the limit experiment. This equivalence follows from two key results. The first, previously demonstrated in Adusumilli (2025b), states that the likelihood ratio processes in the original experiment converge uniformly, at all possible ‘attention’ values, to their counterparts in the limit experiment. We restate this result here for completeness. The second result, which is novel to this article, asserts that any sequence of score and allocation processes, $\{x_{n,a}(\cdot),q_{n,a}(\cdot)\}_{a}$, admits a corresponding representation in the limit experiment.
@@ 3.1. Convergence of likelihood ratio processes
In Adusumilli (2025b), this author showed that under Assumption 1,
@@
By itself, (3.3) does not reference a specific allocation process. While it describes convergence of likelihood ratio processes across the space of all possible ‘attention’ values $(\gamma_{1},\gamma_{0})$, it does not immediately guarantee convergence at all possible time points $t$. Indeed, the relationship between the observed attention values $(\gamma_{1},\gamma_{0})$ and time $t$ is determined by the specific allocation process being employed.
To demonstrate that a risk function in the finite-sample experiment admits a corresponding representation or lower bound in the limit experiment, we need to go further and match the joint distribution of the asymptotically sufficient statistics $\{x_{n,a}(\cdot),q_{n,a}(\cdot)\}_{a}$ as well. Combining this joint convergence with (3.3) would yield uniform convergence of likelihood ratios across all time points, and thereby imply asymptotic equivalence between the actual and limit experiments in the sense of Le Cam (1986). The following key result establishes this joint convergence.
@@ 3.2. The main result
This ensures that $\{z_{n,a}(\cdot),q_{n,a}(\cdot)\}_{a}$ converges to some weak limit $\{z_{a}(\cdot),q_{a}(\cdot)\}_{a}$. Moreover, the measurability of the events $\{q_{n,1}(t)\leq\gamma_{1},q_{n,0}(t)\leq\gamma_{0}\}$ with respect to $\mathcal{G}_{n,\gamma_{0},\gamma_{1}}:=\mathcal{F}_{n,\gamma_{1}}^{(1)}\lor\mathcal{F}_{n,\gamma_{0}}^{(0)}\lor\sigma(U)$ suggests that $q_{a}(\cdot)$ can be constructed to inherit the appropriate adaptedness properties required by Definition 1. The construction requires some care and utilizes some results from the theory of stable convergence (Häusler and Luschgy, 2015); it is perhaps the most intricate part of the proof. Setting $x_{n,a}(\cdot):=z_{n,a}(q_{n,a}(\cdot))$ and $x_{a}(\cdot):=z_{a}(q_{a}(\cdot))$ then gives the desired result.
@@ 4.1. Bayes risk
Let $m_{\bm{\theta}}(\cdot)$ denote a given prior over $\bm{\theta}$. This induces a local prior, $m(\cdot)$, over $\bm{h}$ through the transformation $\bm{h}=\sqrt{n}(\bm{\theta}-\bm{\theta}_{0})$. We consider an asymptotic regime wherein $m(\cdot)$ is assumed to be independent of $n$. The influence of the prior thus remains asymptotically non-negligible and the Bernstein-von Mises theorem does not apply. As discussed in Adusumilli (2025b), local priors offer a better approximation to finite-sample behavior because their influence does not diminish with sample size.
@@ 5. Application 2: Equivalence of in-sample regret
Let $\mu_{n,a}(\bm{h}):=\mathbb{E}_{n,\bm{h}}[Y_{i}^{(a)}]$ denote the average reward corresponding to treatment $a$ when the local parameter is $\bm{h}$. Following Hirano and Porter (2023) and Adusumilli (2025b), the reference parameter is chosen such that $\mu_{n,a}(\bm{0})=0$. The frequentist regret of any $\pi_{n}\equiv\{\pi_{n,j}\}_{j}$ is given by
@@
Previous work by this author (Adusumilli, 2025b) characterized the optimal Bayes and minimax risks in this setting, and showed that they can be attained by a sequence of policy rules that depend on just $\left\{x_{a}(t),q_{a}(t)\right\}_{a}$, i.e., the past values of these variables are not relevant. Theorem 3 is more general, in that it applies to arbitrary sequences of policy rules, but it makes use of a larger information set that includes the entire sample paths of $\{x_{a}(s);s\leq t\}_{a}$ until time $t$.
@@ 6.1. E-processes in multi-treatment adaptive experiments
Definition 2 generalizes the usual notion of an e-process to a multi-indexed super-martingale, where the indices are the treatment allocation proportions. In this framework, the value of $\varepsilon_{n}(\cdot)$ depends on the trajectory of the empirical allocation process, and the super-martingale property holds across all possible empirical allocation processes, i.e., all possible adaptive experiments. A central feature of the e-process is that it is algorithm-free: it remains a valid supermartingale at any time point of any adaptive experiment.
@@ 6.2. GRO, mGRO and REGROW
Following Ramdas et al. (2023) and Grünwald et al. (2024), a common approach to evaluating e-processes in stopping-time experiments is through the Growth Rate Optimality (GRO) criterion. GRO assesses the quality of an e-process $\varepsilon_{n}(\cdot)$ based on its expected log-growth under an alternative hypothesis $\bm{h}\in\mathcal{H}_{1}$. We extend this criterion to the multi-treatment adaptive setting by defining the GRO score
@@
An alternative criterion, following Grünwald et al. (2024), is the REGROW (RElative GRowth Optimality in Worst case) score:
@@ 6.3. Local asymptotics and e-processes in the limit experiment
We then define an equivalent notion of an e-process in the limit experiment. Recall that the limit experiment is characterized by Gaussian process signals $z_{a}(q):=I_{a}^{1/2}h^{(a)}q+W_{a}(q)$. The e-process in the limit experiment is a $\mathcal{G}_{q_{1},q_{0}}$-adapted stochastic process designed for testing $H_{0}:\bm{h}\in\mathcal{H}_{0}$ versus $H_{1}:\bm{h}\in\mathcal{H}_{1}$. The formal definition is given below. First, a bit of terminology: let $\bm{\mathcal{Q}}$ denote the collection of all allocation processes in the limit experiment that are the weak limits of some possible sequence of empirical allocation processes $\{q_{n,a}(\cdot)\}_{a}$.
@@
Relative to Definition 2, the above definition restricts the set of allocation processes to $\bm{\mathcal{Q}}$. The rationale behind this is mainly technical: it allows us to avoid dealing with $\{q_{a}(\cdot)\}_{a}$ that cannot be generated by any sequence of experimental protocols in the actual experiment. Whether $\bm{\mathcal{Q}}$, in fact, includes all possible allocation processes is currently unknown (to this author).
@@ 6.5. Applying the representation theorems
In what follows, we simplify matters by assuming that the null hypothesis $\mathcal{H}_{0}$ is a singleton, consisting solely of the reference parameter $\bm{\theta}_{0}$. [^8] When there is only adaptive stopping, compound nulls are typically addressed in anytime-valid inference using the method of reverse information projection (see Ramdas et al., 2023, for a survey). Extending this approach to the adaptive sampling setting is more involved and left for future work. See, however, Appendix E for the simpler case of testing parameters corresponding to a single treatment arm. Theorem 4 and Corollaries 3, 4 establish asymptotic upper bounds on the GRO, mGRO and REGROW criteria. These bounds are obtained by optimizing the respective criteria within the limit experiment.
@@ 6.5.2. REGROW optimality
Grünwald et al. (2024) show that the REGROW-optimal e-value, $\bar{\varepsilon}_{q_{1}(t),q_{0}(t)}^{*}$, coincides with the mGRO-optimal e-value for a specific, least-favorable weighting function $w_{q_{1}(t),q_{0}(t)}^{*}(\cdot)$. The authors also show that this weighting function is obtained as the solution to
@@
Grünwald et al. (2024) further demonstrate that the optimized value of the objective in (6.5) provides an upper bound on the REGROW criterion $\mathcal{R}\left(\varepsilon;\{q_{a}(t)\}_{a}\right)$:
@@
Crucially, the least-favorable distribution $w_{(\cdot)}^{*}$ depends on both the chosen allocation process and the structure of the alternative hypothesis set $\mathcal{H}_{1}$. It is essential that $\mathcal{H}_{1}$ be compact—without it, the channel capacity (and hence the REGROW value) becomes infinite. Because of the dependence on the allocation process, a globally optimal REGROW e-process does not exist.
As an alternative, one could fix a specific allocation process $\{q_{a}(t)\}_{a}$ and an alternative region $\mathcal{H}_{1}$, from which an optimal weighting function $w_{(\cdot)}^{*}$ can be derived, e.g., by employing the Blahut-Arimoto algorithm (see, Arimoto, 1972; Blahut, 1972). The resulting e-process is thus locally REGROW optimal relative to the chosen design and alternative set (it would also be globally mGRO optimal relative to the least favorable distribution $w^{*}(\cdot)$). Appendix E provides an example of such a locally REGROW optimal e-process, constructed to be optimal against fixed values of $\{q_{a}\}_{a}$.
@@ 7. Conclusion
This article introduces a continuous-time formalism for analyzing fully adaptive experiments. The formalism is based on the notion of allocation processes, also introduced in this article. We show that any empirical allocation process, as induced by some policy rule, converges weakly to a corresponding allocation process in a limit experiment governed by Gaussian diffusions. The various applications, ranging from point-estimation to anytime-valid inference illustrate the utility of this framework.
Beyond these applications, the continuous-time formulation offers a powerful tool for addressing design problems in adaptive experimentation. Though not based on the results reported in this article, prior work by this author has applied the continuous-time framework to derive optimal algorithms for bandit experiments and costly sampling problems (Adusumilli 2025b, a). The current results provide an easy-to-use and generalizable template for transferring optimal designs from the limit experiment back to the finite-sample setting. Looking ahead, we expect this approach to be broadly useful in addressing a range of open questions in adaptive experimentation—e.g., in deriving optimal strategies for best-arm identification with multiple treatments, or in designing adaptive experiments in strategic environments involving interactions between an experimenter and a regulator.
@@ A.1. Proof of Lemma 1
Fix any $t_{1},t_{2}\in[0,1]$ such that $t_{2}\geq t_{1}$, and define $\tau_{1}:=q_{1}(t_{1}),\tau_{2}:=q_{1}(t_{2})$. Also, for each $\gamma_{1}\geq 0$, define $\mathcal{H}_{\gamma_{1}}:=\mathcal{G}_{\gamma_{1},1}$. By the definition of the allocation process, the event $\{q_{1}(t)\leq\gamma_{1}\}$ is $\mathcal{H}_{\gamma_{1}}$ measurable for all $\gamma_{1}$, under any given $t$. Hence, $\tau_{1},\tau_{2}$ are both $\{\mathcal{H}_{\gamma_{1}}\}_{\gamma_{1}\geq 0}$-adapted stopping times. It is easily verified from the definition of $\mathcal{G}_{\gamma_{1},\gamma_{0}}$ that $z_{1}(\gamma_{1})$ is a Wiener process with respect to $\mathcal{H}_{\gamma_{1}}$. Since $\tau_{2}\geq\tau_{1}$ almost surely (a.s.,) due to the almost sure monotonicity of $q_{1}(\cdot)$, it follows by Doob’s optional sampling theorem that $\mathbb{E}\left[z(\tau_{2})|\mathcal{H}_{\tau_{1}}\right]=z(\tau_{1})$ a.s. In other words, $\mathbb{E}\left[z_{1}(q_{1}(t_{2}))|\mathcal{H}_{q_{1}(t_{1})}\right]=z_{1}(q_{1}(t_{1}))$ a.s., i.e., $\mathbb{E}\left[x_{1}(t_{2})|\mathcal{H}_{q_{1}(t_{1})}\right]=x_{1}(t_{1})$. Since $\mathcal{I}_{t}\equiv\mathcal{G}_{q_{1}(t),q_{0}(t)}\subseteq\mathcal{G}_{q_{1}(t),1}\equiv\mathcal{H}_{q_{1}(t)}$ for any $t$ and $x_{1}(t)$ is $\mathcal{I}_{t}$-adapted, the tower property of conditional expectations then implies $\mathbb{E}\left[x_{1}(t_{2})|\mathcal{I}_{t_{1}}\right]=x_{1}(t_{1})$ a.s. But $t_{1},t_{2}$ were arbitrary, so the claim follows.